My dear friends, if anyone can help I would be so grateful. Solve by substitution x+5y=0
-2x-4y=6
Thanks.

Answer by checkley71(8403)

X+5Y=0 OR X=-5Y NOW REPLACE X IN THE OTHER EQUATION BY(-5Y) & SOLVE FOR Y
-2(-5Y)-4Y=6
10Y-4Y=6
6Y=6
Y=6/6
Y=1 ANSWER.
X=-5*1
X=-5 ANSWER.
PROOF
-2*-5-4*1=6
10-4=6
6=6